The Ten Commandments May 27, 2007

Physics 213 General Physics I

SIR ISAAC N EWT ON′S T EN COMMAN DMEN T S

as Recorded by His Servant, Dr. Itai Seggev

1. Thou shalt draw thy free-body diagram for each object in the problem.

2. Thou shalt pick a sensible coordinate system.

3. Thou shalt write down Newton’s second law in components and, if necessary, the torque equation (∑ τ = 0 or ∑ τ = Iα) for each object in the problem.

4. Remember thy sign conventions to keep your equations consistent.

5. Honor thy Professor and thy Grader.

6. If thou hast more unknowns than equations, thou shalt look for additional relationships from kinematics or geometry.

7. Thou shalt identify the desired quantity and solve for it symbolically.

8. Thou shalt plug in the numbers at the end.

9. Thou shalt include units with every number.

10. Thou shalt gaze upon thy final answer to determine if it is at all reasonable. If it is not, thou shalt repeat this process.

Copyright Notice

This document is copyrighted © 2005-7 by Itai Seggev. It may, however, be freely redistributed or used in derivative works for noncommercial purposes and with proper attribution under the terms of the Creative Commons Attribution-NonCommercial-ShareAlike License, Version 3.0. This license should be held to apply equally to all forms of this document, whether print or electronic, including but not limited to any current or future file format. However, the requirements (i) that the full copyright notice be included; (ii) that the disclaimer of warranty be included; and (iii) that the license text or its URI be included; as detailed in sections 4(a) and 4(d) of the license, are hereby waived by the author provided all of the following conditions are met:

1. the copies are distributed exclusively to students enrolled in a course—which is taught by the licensee—at a secondary or post-secondary school;

2. the copies are distributed in hard copy or binary-only form (e.g., PDF); if distributed in binary form via a computer network, the URI of the copy must be restricted to students enrolled in the course;

3. the author is credited in the following or some substantially similar form: “Written by Itai Seggev” for copies of the Work and “Based on a Work by Itai Seggev” for derivatives of the Work.

For details of the license see http://www.creativecommons.org/licenses/by-nc-sa/3.0/.